OpenStudy (anonymous):

Simplify 12x^3 - 8x^2y - 16x/4x =

thomaster (thomaster):

\[12x^3-8x^2y-\frac{ 16x }{ 4x }\] \[12x^3-8x^2y-\frac{ 16x*x }{ 4 }\] \[12x^3-8x^2y-\frac{ 4x^2 }{ 1 }\] \[12x^3-8x^2y-4x^2\]

OpenStudy (anonymous):

love u

OpenStudy (anonymous):

i think it should be simplified more

thomaster (thomaster):

It don't see how it can be simplified further

OpenStudy (anonymous):

3x4+6x3+x2 3x2-2xy-4 3x2+2x+1 None of the above

thomaster (thomaster):

oh you can divide it by 4 so you'll get \[3x^3-2x^2y-x^2\]So i guess it's none of the above

OpenStudy (anonymous):

it is wrong

OpenStudy (anonymous):

\[4x ^{3} - 2x ^{2}y - 1\]

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

theomaster, the second step is wrong. You can't do that, its dividing not multiplying

thomaster (thomaster):

I see :P something went wrong there. 16x/4x=4 But the answer would still be none of the above

OpenStudy (anonymous):

yes